OpenStudy (anonymous):

8x squared +64x factor the expression

OpenStudy (anonymous):

(8x) ^ 2 + 64 x = 64x^2 + 64x = 64x(x + 1) or if you mean 8x^2 + 64x = 8x(8x + 8)

OpenStudy (anonymous):

8x(x+8)

OpenStudy (anonymous):

^Right.